Fickel Company has two manufacturing departments—Assembly and Testing & Packaging. The predetermined overhead rates in Assembly and Testing & Packaging are $25.00 per direct labor-hour and $21.00 per direct labor-hour, respectively. The company’s direct labor wage rate is $27.00 per hour. The following information pertains to Job N-60:

Assembly Testing & Packaging
Direct materials $ 405 $ 51
Direct labor $ 297 $ 54

Required:

1. What is the total manufacturing cost assigned to Job N-60? (Do not round intermediate calculations.)

2. If Job N-60 consists of 10 units, what is the unit product cost for this job? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1
Direct labor hours:
Assembly 11 =297/27
Testing & Packaging 2 =54/27
1
Direct materials 456 =405+51
Direct labor 351 =297+54
Overhead applied:
Assembly 275 =11*25
Testing & Packaging 42 =2*21
Total manufacturing cost 1124
2
Total manufacturing cost 1124
Divide by Units 10
Unit product cost 112.40 per unit